Determine net radiant-heat exchange between the closed ends of a hollow cylinder 12-in. in diameter and 4-in. long. The ends are of Cr-Ni alloy with one end heated electrically to a temperature of 1900 °F and the other end maintained at 125 °F. The side wall of the cylinder is substantially nonconducting and reradiating.
1
Expert's answer
2021-10-29T02:32:48-0400

Heat transfer can be expressed with "Fourier's Law"

"q = \\frac{k }{s} A dT \n\n\ufeff"

Thermal conducitivity "k = 11.3 \frac{W}{m}" ,s =1


Area "A= \pi DL =\pi \times0.3048\times0.1016 =0.09728 m^2"


"q = 11.3\\times0.09728\\times(1900-125) =1951.19W =1.95kW\n\n\ufeff"